OpenStudy (anonymous):

how do you enter the equation of the circle in standard form with center and radius given. Center (-8, -3), radius = 2

OpenStudy (anonymous):

radius= 2 sqrt 3

OpenStudy (anonymous):

\[(x-h)^2+(y-k)^2=r^2\] with \(h=-8, k=-3, r=2\)

OpenStudy (anonymous):

ok but how do you do it with the sqrt?

OpenStudy (anonymous):

your radius is \(2\sqrt{3}\) right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but the equation needs \(r^2\) so take \(2\sqrt{3}\) and square it (multiply it by itself) clear or no?

OpenStudy (anonymous):

so that would make it 6?

OpenStudy (anonymous):

no you have to square 2 as well

OpenStudy (anonymous):

so 12?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

yw